# RouteLLM: Learning to Route LLMs with Preference Data

### v3

@@ 3 Methodology
3.1 Preference Data
We start by describing how we can obtain the preference data to train the routing function. We primarily use 80k battles from the online Chatbot Arena platform [12]. On this platform, users interact with a chatbot interface and submit prompts of their choice. Upon submission, they receive responses from two anonymous models and vote for a winning model or a tie. The resulting dataset, denoted as $\mathcal{D}_{\text{arena}}=\{(q,a_{i},a_{j},l_{i,j})\mid q\in\mathcal{Q},a_{i},a_{j}\in\mathcal{A},l_{i,j}\in\mathcal{L}\}$, consisting of user queries, answers from two models $M_{i},M_{j}$, and a pairwise comparison label based on human judgment.
A major issue with using the raw Chatbot Arena data is label sparsity. For instance, the percentage of comparison labels between any two models, on average, is less than 0.1%. Therefore, we derive the preference data for training the router as follows: First, we reduce label sparsity by clustering models in $\mathcal{D}_{\text{arena}}$ into 10 different tiers (see Appendix A), using each model’s Elo score on the Chatbot Arena leaderboard [^5] https://leaderboard.lmsys.org and aiming to minimize the variation within each tier via dynamic programming. We choose the models on the first and second tiers to represent strong models $\mathcal{M}_{\text{strong}}$, and the models on the third tier to represent weak models $\mathcal{M}_{\text{weak}}$. While we primarily train on battles across these tiers, we also leverage battles involving other model tiers to regularize our learning methods. Crucially, we omit the actual model responses in $\mathcal{D}_{\text{arena}}$, retaining only the model identities, i.e. $e\sim\mathcal{D}_{\text{pref}}$ is $e={(q,M_{i},M_{j},l_{i,j})}$. The comparison label $l_{i,j}$ still provides insight into the relative capabilities of the LLMs $M_{i}$ and $M_{j}$ on various types and complexity levels of the query $q$.
@@ 3.1.1 Data Augmentation
Even after classifying models into tiers, the human preference signal can still be quite sparse across different model classes. As we discuss in Sec 4.1, we find that this can hinder generalization, especially for parameter-heavy models. Thus we explore two data augmentation methods:
Golden-labeled datasets: We augment our training data with datasets of the form $\mathcal{D}_{\text{gold}}=\{(q,a,l_{g})\mid q\in\mathcal{Q},a\in\mathcal{A},l_{g}\in\mathbb{R}\}$, where a golden label $l_{g}$ is automatically calculated for a model answer $a$, e.g. multiple-choice answers. One example of such a dataset is the MMLU benchmark [17]. We use the validation split of MMLU containing approximately 1500 questions and derive comparison labels $l_{i,j}$ from $l_{g}$ by simply comparing the responses from $M_{i}$ and $M_{j}$, creating the preference dataset $\mathcal{D}_{\text{gold}}$ for augmentation.
LLM-judge-labeled datasets: We explore obtaining preference labels on open-ended purpose chat domains using a LLM judge [33], as it has demonstrated a high correlation with human judgment [16, 20]. Given a collection of user queries, we start by generating responses from both a strong model in $\mathcal{M}_{\text{strong}}$ and a weak model in $\mathcal{M}_{\text{weak}}$ and then produce pairwise comparison labels using GPT-4 as a judge. The primary challenge with this method is the high cost of collecting responses and pairwise comparisons from GPT-4 in large quantities. Fortunately, the Nectar dataset [34] offers a wide variety of queries with corresponding model responses. We significantly reduce our costs by selecting queries with GPT-4 responses (as a representative of $\mathcal{M}_{\text{strong}}$), on which we generate responses from Mixtral-8x7B (as representative of $\mathcal{M}_{\text{weak}}$). Finally, we obtain pairwise comparison labels using the GPT-4 judge. [^6] We employ best practices recommended in [33] to de-bias GPT-4 judgements. Overall, we collect a preference dataset $\mathcal{D}_{\text{judge}}$ of approximately 120K samples with a total cost of around \$700 USD.
@@ 3.2 Routing Approaches
We now detail our methods for learning the win prediction model $P_{\boldsymbol{\theta}}(\text{win}_{\mathcal{M}_{\text{strong}}}|q)$ from preference data $\mathcal{D}_{\text{pref}}$. We denote a sample $(q,M_{i},M_{j},l_{i,j})\sim\mathcal{D}_{\text{pref}}$ as $e=(q,M_{w},M_{l})$, where $M_{w}$ and $M_{l}$ refer to the winning and losing model respectively.
@@
We adopt a Bradley-Terry (BT) model [7] similar to [12]. Given a user query $q$, we compute a weight $\omega_{i}=\gamma^{1+S(q,\hat{q})}$  [^7] We find that exponential scale works best in practice and choose $\gamma=10$ for each query $q_{i}$ in the train set based on its similarity to $q$:
$\mathcal{S}(q,q_{i})=\frac{\epsilon\cdot\epsilon_{i}}{\lVert\epsilon\rVert\lVert\epsilon_{i}\rVert\cdot\max_{1\leq s\leq\mid\mathcal{D}_{\text{pref}}\mid}\frac{\epsilon_{i}\cdot\epsilon_{s}}{\lVert\epsilon_{i}\rVert\lVert\epsilon_{s}\rVert}},$ (8)
where $\epsilon$ denotes a query embedding. We learn BT coefficients $\xi$ (representing 10 model classes) by solving:
$\mathop{\mathrm{argmin}}_{\xi}\sum_{i=1}^{\mid\mathcal{D}_{\text{pref}}\mid}\left[\omega_{i}\cdot\ell\left(l_{i},\frac{1}{1+e^{\xi_{w_{i}}-\xi_{l_{i}}}}\right)\right],$ (9)
where $\ell$ is a binary cross-entropy loss. The resulting BT coefficients allow us to estimate the win probability as: $P(\text{win}_{M_{w}}|q)=\frac{1}{1+e^{\xi_{w}-\xi_{l}}}$. For this router model, there is no training required and the solving is performed at inference time.
@@
Inspired by matrix factorization models [21, 28] in recommendation systems to capture the low-rank structure of user-item interactions, we leverage this approach for training on preference data. The key is to uncover a hidden scoring function $s:\mathcal{M}\times\mathcal{Q}\rightarrow\mathbb{R}$. The score $s(M_{w},q)$ should represent the quality of the model $M_{w}$’s answer to the query $q$, i.e. if a model $M_{w}$ is better than $M_{l}$ on a query $q$, then $s(M_{w},q)>s(M_{l},q)$. We enforce this relationship by modeling the win probability with a BT relationship [7]:
$P(\text{win}_{M_{w}}|q)=\sigma(s({M}_{w},q)-s({M}_{l},q)),$ (10)
which we optimize on preference data. We model the scoring function $s$ as a bilinear function of model and query, and embed the model identity $M$ to a $d_{m}$-dimensional vector ${v}_{m}$ and the query to a $d_{q}$-dimensional vector ${v}_{q}$:
$s({M},q)=w_{2}^{T}(v_{m}\odot(W_{1}^{T}v_{q}+b))$ (11)
Here, $\odot$ represents the Hadamard product, $W_{1}\in\mathbb{R}^{d_{q}\times d_{m}}$ and $b\in\mathbb{R}^{d_{m}}$ is the projection layer to align the dimension of $v_{q}$ with $v_{m}$, and $w_{2}\in\mathbb{R}^{d_{m}}$ is the linear regression layer to produce the final scalar. This method is essentially learning a matrix factorization of the score matrix on the set $\mathcal{Q}\times\mathcal{M}$. We train the model on a 8GB GPU for $\approx$ 10 epochs, using batch size 64 and the Adam optimizer with learning rate $3\times 10^{-4}$ and weight decay $1\times 10^{-5}$.
@@
Here we explore using a standard text classification method with a higher number of parameters compared to previous methods. We use a BERT-base architecture [14], to give a contextualized embedding of the user query, and define win probability as:
$P_{\boldsymbol{\theta}}(\text{win}_{M_{w}}|q)=\sigma(Wh_{\text{CLS}}+b),$ (12)
where ${h}_{\text{CLS}}$ is an embedding corresponding to the special classification token (CLS) summarizing the input query $q$; and $W,b,\sigma$ are parameters and sigmoid activation of a logistic regression head. We perform full-parameter fine-tuning on $\mathcal{D}_{\text{pref}}$. We train the model on 2xL4 24GB GPUs for $\sim 2000$ steps using a batch size of 16, maximum sequence length of 512, learning rate of $1\times 10^{-5}$ and weight decay of 0.01.
@@
We finally expand the capacity of our router by parameterizing it with Llama 3 8B[3]. We use an instruction-following paradigm [31], i.e. we provide as input an instruction prompt containing the user query, and output the win probability in a next-token prediction fashion – instead of using a separate classification head. Notably, we append the comparison labels as additional tokens to the vocabulary, and compute the win probability as softmax over the label classes $\mathcal{L}$. We train the model on 8xA100 80GB GPUs for $\sim 2000$ steps using a batch size of 8, maximum sequence length of 2048, and a learning rate of $1\times 10^{-6}$.
@@ 4 Experiments
Training data: As mentioned in Sec. 3.1, we primarily use the 80K Chatbot Arena for training our models, but hold out 5k samples for validation. We prune all prompt samples shorter than 16 characters, resulting in 65k pairwise comparisons between 64 different models. These consist of conversations from over 100 languages, with the bulk of the conversations (81%) in English, followed by Chinese (3.1%), and Russian (2.2%). We assign models to 10 classes to reduce sparsity of comparison labels. As discussed in Sec. 3.1.1, we further augment our training data with with either: 1) $\mathcal{D}_{\text{gold}}$, golden-labeled data created from the MMLU validation split and 2) $\mathcal{D}_{\text{judge}}$, GPT-4-as-a-judge labeled chat data.
Evaluation benchmarks: We evaluate our routers on three widely-used academic benchmarks: MMLU [17] consisting of 14,042 questions across 57 subjects, MT Bench [33] with 160 open-ended questions using LLM-as-a-judge, and GSM8K [13] with over 1,000 grade school math problems. Additionally, we conduct a cross-contamination check between our evaluation and training datasets, and report uncontaminated results below. We present results on public benchmarks to understand the out-of-domain generalization of routers.
Routers: For both the matrix factorization router and the similarity-weighted ranking router, we use OpenAI’s embedding model text-embedding-3-small to embed the input query. We perform full-parameter finetuning on both BERT and Causal LLM, and use the validation set for model selection. We opt to use gpt-4-1106-preview [23] as a representative model in $\mathcal{M}_{\text{strong}}$, and Mixtral 8x7B [19] as a representative model in $\mathcal{M}_{\text{weak}}$, to concretely evaluate router performance. We also use a random router that routes queries randomly under a cost constraint as a baseline.
@@ 4.1 Results
Training data Method $CPT(50\%)$ $CPT(80\%)$ $APGR$
@@
Table 1: MT Bench results. Note that the score for CPT at 50% (8.8) is 95% that of GPT-4 performance (9.3). Our routers exhibit strong performance on MT Bench when trained on $\mathcal{D}_{\text{arena}}$, with further improvement when the dataset is augmented with $\mathcal{D}_{\text{judge}}$, reducing costs by up to 75% as compared to the random router.
@@
Augmenting the preference data using a GPT-4 judge leads to notable improvements across all routers. The BERT and causal LLM routers now perform much better than the random baseline, with the BERT classifier achieving an APGR improvement of over 50% as compared to random. When trained on this augmented dataset, matrix factorization is the best-performing router as its CPT(80%) is nearly halved, requiring 50% less GPT-4 calls as compared to random.
We also compare the MT Bench performance of our routers against existing routing systems in Appendix E, demonstrating the substantial improvements that our routers achieve over other available systems.
Training data Method $CPT(50\%)$ $CPT(80\%)$ $APGR$
@@
Table 2: 5-shot MMLU results for our routers. Note that the score for CPT at 50% (75) is 92% that of GPT-4 performance (81). Routers trained only on $\mathcal{D}_{\text{arena}}$ perform poorly due to most questions being out-of-distribution, but dataset augmentation with $\mathcal{D}_{\text{gold}}$ is highly effective, leading to significant improvement in router performance even with a small number of samples.
@@
Training data Method $CPT(50\%)$ $CPT(80\%)$ $APGR$
@@
Table 3: 8-shot GSM8K results. Note that the score for CPT at 50% (75) is 87% that of GPT-4 performance (86). Routers trained only on $\mathcal{D}_{\text{arena}}$ again perform poorly due to questions being out-of-distribution, but augmentation with $\mathcal{D}_{\text{judge}}$ substantially improves router performance.
Finally, on GSM8K (Table 3), we observe that similar to MMLU, the performance of all routers trained only on the Arena dataset is close to random. However, training our routers on the dataset augmented with synthetic data from an LLM judge improves performance substantially, with all routers going from an APGR worse than random to an APGR greater than random. When trained on this augmented dataset, the causal LLM classifier performs the best out of all routers, requiring 17% less GPT-4 calls than random to achieve CPT(50%) and CPT(80%).

### v4

@@ 4 Methodology
4.1 Chatbot Arena Data
Our primary source for preference data is the 80k battles from the online Chatbot Arena platform (Chiang et al., 2024), where users submit prompts and receive responses from two anonymous models. After reviewing the responses, users vote for a winner or declare a tie. This generates a dataset, denoted as $\mathcal{D}_{\text{arena}}$, which contains user queries, model responses, and pairwise comparison labels based on human judgment.
A key challenge with the raw Chatbot Arena data is label sparsity. On average, the percentage of comparison labels between any two models is less than 0.1%. To address this, we derive the preference data by clustering the models in the data into 10 tiers (see Appendix A) based on their scores on the Chatbot Arena leaderboard [^3] https://leaderboard.lmsys.org, and minimize intra-tier variation using dynamic programming. We choose models in the top two tiers to represent the $\mathcal{M}_{\text{strong}}$ class, and models in the third tier represent the $\mathcal{M}_{\text{weak}}$ class. Crucially, we exclude model responses and retain only the winner identities in training. The resulting dataset is defined as $\mathcal{D}_{\text{arena}}=\{(q,l_{s,w})\mid q\in\mathcal{Q},l_{s,w}\in L\}$.
@@ 4.1.1 Data Augmentation
Despite classifying models into tiers, the human preference signal remains sparse across different model classes. As discussed in Sec 5.1, this sparsity can limit generalization, particularly for parameter-heavy router models. To address this, we explore two data augmentation methods:
Golden-labeled datasets: We augment our training data with labeled datasets of the form $\mathcal{D}_{\text{gold}}=\{(q,l_{g},l_{s,w})\mid q\in\mathcal{Q},l_{g}\in\mathbb{R},l_{s,w}\in L\}$, where a golden label $l_{g}$ is the known correct answer, e.g. in multiple-choice questions. Specifically, we use the validation split of the MMLU multiple choice benchmark (Hendrycks et al., 2020) containing approximately 1500 questions and derive comparison labels $l_{s,w}$ simply by comparing the responses from $M_{s}$ and $M_{w}$ to the golden label.
LLM-judge-labeled datasets: We explore obtaining preference labels on open-ended purpose chat domains using a LLM judge (Zheng et al., 2023), as it has demonstrated a high correlation with human judgment (Dubois et al., 2024; Jiang et al., 2023). Given a collection of user queries, we start by generating responses from both a strong model $M_{s}\in\mathcal{M}_{\text{strong}}$ and a weak model $M_{w}\in\mathcal{M}_{\text{weak}}$, then producing pairwise comparison labels using GPT-4 as a judge. The primary challenge with this method is the high cost of collecting responses and pairwise comparisons from GPT-4 in large quantities. Fortunately, the Nectar dataset (Zhu et al., 2023) offers a wide variety of queries with corresponding model responses. We significantly reduce our costs by selecting queries with GPT-4 responses (as $M_{s}$), on which we generate responses from Mixtral-8x7B (as $M_{w}$). Finally, we obtain pairwise comparison labels using the GPT-4 judge. [^4] We employ best practices recommended in (Zheng et al., 2023) to de-bias GPT-4 judgements. Overall, we collect a preference dataset $\mathcal{D}_{\text{judge}}$ of approximately 120K samples costing around \$700 USD in total.
@@ 4.2 Routing Approaches
We now discuss several methods to define the win prediction model $P_{\bm{\theta}}(\text{win}_{s}|q)$ introduced in Eq 1.
@@
We adopt a Bradley-Terry (BT) model (Bradley & Terry, 1952) similar to Chiang et al. (2024). Given a user query $q$, we compute its cosine similarity to each query $q^{\prime}$ in the train set, scaled according to the maximum cosine similarity for $q^{\prime}$ in the dataset:
$\mathcal{S}(q,q^{\prime})=\frac{\epsilon\cdot\epsilon^{\prime}}{\lVert\epsilon\rVert\lVert\epsilon^{\prime}\rVert\cdot\max_{1\leq s\leq\mid\mathcal{D}_{\text{pref}}\mid}\frac{\epsilon^{\prime}\cdot\epsilon_{s}}{\lVert\epsilon^{\prime}\rVert\lVert\epsilon_{s}\rVert}},$ (9)
where $\epsilon$ and $\epsilon^{\prime}$ denote text embeddings for $q$ and $q^{\prime}$ respectively. This similarity score is used to compute a weight scalar for each training query $\omega^{\prime}=\gamma^{1+S(q,q^{\prime})}$ [^5] We find that exponential scale works best in practice and choose $\gamma=10$. We learn BT coefficients $\xi_{s}$, $\xi_{w}$ for the strong and weak models by solving:
$\mathop{\mathrm{argmin}}_{\xi_{s},\xi_{w}}\sum_{(q,l_{s,w})\in\mathcal{D}_{\text{pref}}}\left[\omega^{\prime}\cdot\ell\left(l_{s,w},\frac{1}{1+e^{\xi_{s}-\xi_{w}}}\right)\right],$ (10)
where $\ell$ is a binary cross-entropy loss. These coefficients allow us to estimate the win probability as: $P_{\bm{\theta}}(\text{win}_{s}|q)=\frac{1}{1+e^{\xi_{s}-\xi_{w}}}$. In this approach, no training is required—solving is performed at inference time.
@@
Drawing inspiration from matrix factorization models used in recommendation systems to capture low-rank structures in user-item interactions (Koren et al., 2009; Töscher et al., 2009), we apply this approach to learn from preference data. The goal is to uncover a hidden scoring function $\delta:\mathcal{M}\times\mathcal{Q}\rightarrow\mathbb{R}$, where $\delta(M,q)$ represents the quality of the model $M$’s response to query $q$. If $M$ is better than $M^{\prime}$ on a query $q$, then $\delta(M,q)>\delta(M^{\prime},q)$. We enforce this by modeling the win probability using a sigmoid function $\sigma$:
$P_{\bm{\theta}}(\text{win}_{s}|q)=\sigma\left(\delta(M,q)-\delta(M^{\prime},q)\right),$ (11)
which we optimize on the preference data. The scoring function $\delta$ is modelled as a bilinear function of the model and query embeddings. We embed the model $M$ into a $d_{m}$-dimensional vector ${v}_{m}$, and the query $q$ into a $d_{q}$-dimensional vector ${v}_{q}$:
$\delta({M},q)=w_{2}^{T}(v_{m}\odot(W_{1}^{T}v_{q}+b)),$ (12)
where $\odot$ denotes the Hadamard product. $W_{1}\in\mathbb{R}^{d_{q}\times d_{m}}$ and $b\in\mathbb{R}^{d_{m}}$ are parameters of a projection layer to align the dimension of $v_{q}$ with $v_{m}$. $w_{2}\in\mathbb{R}^{d_{m}}$ is the linear regression layer to produce the final scalar. This method is essentially learning a matrix factorization of the score matrix on the set $\mathcal{Q}\times\mathcal{M}$. We train the model on a 8GB GPU for $\approx$ 10 epochs, using batch size 64 and the Adam optimizer (Kingma & Ba, 2017) with learning rate $3\times 10^{-4}$ and weight decay $1\times 10^{-5}$.
@@
We explore using a standard text classification method with a higher number of parameters compared to previous methods. We use a $\text{BERT}_{\text{BASE}}$ architecture (Devlin et al., 2018), to give a contextualized embedding of the user query, and define win probability as:
$P_{\bm{\theta}}(\text{win}_{s}|q)=\sigma(Wh_{\text{CLS}}+b),$ (13)
where ${h}_{\text{CLS}}$ is an embedding corresponding to the special classification token (CLS) summarizing the input query $q$. $W$ and $b$ are parameters of the logistic regression head, while $\sigma$ is the sigmoid function. We perform full-parameter fine-tuning on $\mathcal{D}_{\text{pref}}$. We train the model on 2xL4 24GB GPUs for $\sim 2000$ steps using a batch size of 16, maximum sequence length of 512, learning rate of $1\times 10^{-5}$ and a weight decay of 0.01.
@@
We finally expand the capacity of our router by parameterizing it with Llama 3 8B (AI@Meta, 2024b). We use an instruction-following paradigm (Wei et al., 2021), i.e. we provide as input an instruction prompt containing the user query and output the win probability in a next-token prediction fashion, instead of using a separate classification head. Notably, we append the comparison labels as additional tokens to the vocabulary, and compute the win probability as a softmax over the label classes $\mathcal{L}$. We train the model on 8xA100 80GB GPUs for $\sim 2000$ steps using a batch size of 8, maximum sequence length of 2048, and a learning rate of $1\times 10^{-6}$.
@@ 5 Experiments
Training data: As mentioned in Sec. 4.1, we primarily use the 80K Chatbot Arena data $\mathcal{D}_{\text{arena}}$ for training our models, but hold out 5k samples for validation. We prune all prompt samples shorter than 16 characters, resulting in 65k pairwise comparisons between 64 different models. These consist of conversations from over 100 languages, with the bulk of the conversations (81%) in English, followed by Chinese (3.1%), and Russian (2.2%). We assign models to 10 classes to reduce sparsity of comparison labels. As discussed in Sec. 4.1.1, we further augment our training data with with either: 1) $\mathcal{D}_{\text{gold}}$, golden-labeled data created from the MMLU validation split, or 2) $\mathcal{D}_{\text{judge}}$, GPT-4-as-a-judge labeled data.
Evaluation benchmarks: We evaluate our routers on three widely-used academic benchmarks: MMLU (Hendrycks et al., 2020) consisting of 14,042 questions across 57 subjects, MT Bench (Zheng et al., 2023) with 160 open-ended questions using LLM-as-a-judge, and GSM8K (Cobbe et al., 2021) with over 1,000 grade school math problems. Additionally, we conduct a cross-contamination check between our evaluation and training datasets, and report uncontaminated results below. We present results on public benchmarks to understand the out-of-domain generalization of our routers.
Routers: For both the matrix factorization router and the SW ranking router, we use OpenAI’s embedding model text-embedding-3-small to embed the input query. We perform full-parameter finetuning on both BERT and Causal LLM, and use the validation set for model selection. We opt to use gpt-4-1106-preview (OpenAI, 2023) as $M_{s}\in\mathcal{M}_{\text{strong}}$ and Mixtral 8x7B (Jiang et al., 2024) as $M_{w}\in\mathcal{M}_{\text{weak}}$ to concretely evaluate router performance. We use a random router that routes queries randomly under a cost constraint as the baseline.
@@ 5.1 Results
Training data Method $CPT(50\%)$ $CPT(80\%)$ $APGR$ Improvement
@@
Table 1: MT Bench results. Note that the MT Bench score at CPT(50%), 8.8, is 95% that of GPT-4’s score (9.3). Our routers exhibit strong performance on MT Bench when trained on $\mathcal{D}_{\text{arena}}$, with further improvement when the dataset is augmented with $\mathcal{D}_{\text{judge}}$, reducing costs by up to 75% as compared to the random router.
@@
Augmenting the preference data using a GPT-4 judge leads to notable improvements across all routers. The BERT and causal LLM routers now perform much better than the random baseline, with the BERT classifier achieving an APGR improvement of over 50% as compared to random. When trained on this augmented dataset, matrix factorization is the best-performing router with its CPT(80%) nearly halved and requiring 50% less GPT-4 calls as compared to random.
We also compare the MT Bench performance of our routers against existing commercial routing systems in Appendix E, demonstrating how our routers achieve substantial improvements over other available systems.
Training data Method $CPT(50\%)$ $CPT(80\%)$ $APGR$ Improvement
@@
Table 2: 5-shot MMLU results for our routers. Note that the MMLU score at CPT(50%), 75, is 92% that of GPT-4’s score (81). Routers trained only on $\mathcal{D}_{\text{arena}}$ perform poorly due to most questions being out-of-distribution, but dataset augmentation with $\mathcal{D}_{\text{gold}}$ is highly effective, leading to significant improvement in router performance even with a small number of samples.
@@
Training data Method $CPT(50\%)$ $CPT(80\%)$ $APGR$ Improvement
@@
Table 3: 8-shot GSM8K results. Note that the GSM8K score at CPT(50%), 75, is 87% that of GPT-4’s score (86). Routers trained only on $\mathcal{D}_{\text{arena}}$ again perform poorly due to questions being out-of-distribution, but augmentation with $\mathcal{D}_{\text{judge}}$ substantially improves router performance.
Finally, on GSM8K (Table 3), we observe that similar to MMLU, the performance of all routers trained only on the Arena dataset is close to random. However, training our routers on the dataset augmented with synthetic data from an LLM judge substantially improves performance, with all routers going from an APGR worse than random to an APGR greater than random. When trained on this augmented dataset, the causal LLM classifier performs the best out of all routers, requiring 17% less GPT-4 calls than random to achieve CPT(50%) and CPT(80%).
5.2 Adaptability across models
We picked gpt-4-1106-preview and Mixtral 8x7B as $M_{s}$ and $M_{w}$ respectively for the above experiments. However, to demonstrate the adaptability of our routers to new LLMs, we report in Table 4 the performance of our routers on MT Bench when they are used to route between two new model pairs: (1) $M_{s}=$ Claude 3 Opus, $M_{w}=$ Claude 3 Sonnet (Anthropic, 2024) and (2) $M_{s}=$ Llama 3.1 70B, $M_{w}=$ Llama 3.1 8B (AI@Meta, 2024a). Importantly, we use the same routers as before without any retraining, and only replace the strong and weak model routed to. These LLMs are also not present in the training data.
Model Pair Method $CPT(50\%)$ $CPT(80\%)$ $APGR$ Improvement
($M_{s}$ / $M_{w}$)
Claude 3 Opus / Random (95% CI) 49.89 ($\pm$3)72.27 ($\pm$4)0.493 ($\pm$0.033) (+0%)
Claude 3 Sonnet BERT 34.85% 39.04% 0.682 (+38.3%)
Causal LLM 28.12% 50.00% 0.656 (+33.1%)
Matrix Factorization 31.86% 36.43% 0.762 (+54.6%)
SW Ranking 23.27% 51.85% 0.772 (+56.6%)
Llama 3.1 70B / Random (95% CI) 47.52($\pm$3)76.26($\pm$2)0.512 ($\pm$0.017) (+0%)
Llama 3.1 8B BERT 30.15% 38.91% 0.673 (+31.4%)
Causal LLM 34.05% 45.96% 0.689 (+34.6%)
Matrix Factorization 25.83% 37.30% 0.738 (+44.1%)
SW Ranking 21.18% 29.39% 0.767 (+49.8%)
Table 4: MT Bench results for our routers when used to route between different model pairs. We use the exact same routers as before trained on $\mathcal{D}_{\text{arena}}+\mathcal{D}_{\text{judge}}$. Our routers generalize very well across different model pairs without any retraining.
We observe strong results across all existing routers on MT Bench even when the model pair is replaced, with performance comparable to that of the original model pair. The results continue to be significantly stronger than random, with our best routers requiring approximately half the GPT-4 calls of the random router to achieve CPT(80%) when routing between both the Claude 3 and Llama 3.1 family of models. These results suggest that our routers have learned common characteristics of queries that allow them to distinguish between strong and weak models, generalizing to new models at inference time without additional training.